OpenStudy (anonymous):

Describe the balance of forces that hold together the nuclei of atoms. What happens when these forces get out of balance? **Not sure how to explain this! Thank you:)

OpenStudy (michele_laino):

the force which is acting among protons and neutrons, is the so called strong nuclear force. Such force can be explained, introducing a particle mediator of that nuclear force. Such mediator particle, is the pi meson (pion) , whose mass is about 140 MeV. In other words, as electromagnetic force is due to the mediator which is the photon, in the same way the nuclear force is due to an exchange, among the nucleons, of a pion, or meson pi, in fact the nuclear force is described inside a theory which is calle OPEP namely One Pion Exchange Potential

OpenStudy (anonymous):

ohhh okay! and so when they get out of balance, what happens? :/

OpenStudy (michele_laino):

when a nucleon, for example a proton, is out of the range of the nuclear force, then on it will act the electrostatic force, so it will repelled by other protons. When a neutron is out of the range of a nuclear force, then that nucleus will become instable

OpenStudy (michele_laino):

the nuclear force, acting among nucleons, is very attractive, nevertheless its range it is very short, that range is about 10^-14 cm

OpenStudy (anonymous):

ohhh okay! thank you!! :D

OpenStudy (anonymous):

and so this problem is complete? :O

OpenStudy (michele_laino):

yes! I think so! If you want I can give you a simple computation, of the range of the strong nuclear force

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

keep in mind that the mediator particle, namely the pion or pi meson , was introduced by the famous physicist Hideki Yukawa (Japan) in the past 1934

OpenStudy (michele_laino):

Now, we can consider the pion a so called "virtual particle". It is virtual not, because it doesn't exist, it is called virtual because its existence violates the Heisenberg's Uncertainty Principle

OpenStudy (anonymous):

Oh! Okay:)

OpenStudy (michele_laino):

so, we can write this equation: \[\Large \Delta E\Delta t \leqslant \hbar \] it is the violation of the Heisenberg's Uncertainty Principle, as you can see

OpenStudy (michele_laino):

Now, we can replace \Delat E, with the mass of the pion, namely: \[\Large \Delta E = {M_\pi }{c^2} = 140MeV\]

OpenStudy (anonymous):

oohhh okay!

OpenStudy (michele_laino):

then we can solve that equation for \Delta t, being \Delta t the existence time of the pion, so we get: \[\Large \Delta t \simeq \frac{\hbar }{{140}}\]

OpenStudy (michele_laino):

Now, since the velocity of the pion is close to the light speed, then we can find the spoace traveled by our pion, using this formula: \[\Large d = c\Delta t = \frac{{\hbar c}}{{140}} = \frac{{197}}{{140}} = 1.4fm\] being \[\Large \hbar c = 197MeV \cdot fm\] and \[\Large 1fm = {10^{ - 13}}cm\] fm=fermi

OpenStudy (anonymous):

ohh okay!! :O

OpenStudy (anonymous):

thank you!!

OpenStudy (michele_laino):

thank you!!